Grande03:

Assuming that Hamid’s revenue is found by multiplying the book price by the number of copies sold, create an equation (in standard form) to model his revenue. Remember that he sold 35 copies in the first week for $10 per copy. After that, every 50-cent increase, x, in the price per copy resulted in the average number of books he sold decreasing by 2 per week.

lilbaddieeeeee:

y = -x^2 − 2.5x + 350 Step-by-step explanation: y = (10 + 0.5x)(35 − 2x) y = 350 − 20x + 17.5x − x^2 y = 350 − 2.5x − x^2

lilbaddieeeeee:

dose this help?

Grande03:

im trying it hold on sorry

Grande03:

yes thank you